OpenStudy (anonymous):

Write a simplified polynomial expression to represent the area of the square tile, shown below

OpenStudy (anonymous):



OpenStudy (anonymous):

please help

jimthompson5910 (jim_thompson5910):

This is a square with side length x-3 so use the formula A = s^2 where s is the side length

OpenStudy (anonymous):

is it the x-3?

jimthompson5910 (jim_thompson5910):

yes so A = s^2 A = (x-3)^2 A = ???

OpenStudy (anonymous):

x-3^2